# TalkingEyes: Pluralistic Speech-Driven 3D Eye Gaze Animation

### v1

@@ Abstract
Although significant progress has been made in the field of speech-driven 3D facial animation recently, the speech-driven animation of an indispensable facial component, eye gaze, has been overlooked by recent research. This is primarily due to the weak correlation between speech and eye gaze, as well as the scarcity of audio-gaze data, making it very challenging to generate 3D eye gaze motion from speech alone. In this paper, we propose a novel data-driven method which can generate diverse 3D eye gaze motions in harmony with the speech. To achieve this, we firstly construct an audio-gaze dataset that contains about 14 hours of audio-mesh sequences featuring high-quality eye gaze motion, head motion and facial motion simultaneously. The motion data is acquired by performing lightweight eye gaze fitting and face reconstruction on videos from existing audio-visual datasets. We then tailor a novel speech-to-motion translation framework in which the head motions and eye gaze motions are jointly generated from speech but are modeled in two separate latent spaces. This design stems from the physiological knowledge that the rotation range of eyeballs is less than that of head. Through mapping the speech embedding into the two latent spaces, the difficulty in modeling the weak correlation between speech and non-verbal motion is thus attenuated. Finally, our TalkingEyes, integrated with a speech-driven 3D facial motion generator, can synthesize eye gaze motion, eye blinks, head motion and facial motion collectively from speech. Extensive quantitative and qualitative evaluations demonstrate the superiority of the proposed method in generating diverse and natural 3D eye gaze motions from speech. The project page of this paper is: https://lkjkjoiuiu.github.io/TalkingEyes_Home/.
Index Terms:
Eye Gaze, Speech-driven, 3D Facial Animation
@@ I Introduction
The computer graphics and vision community has witnessed a significant progress in the field of speech-driven 3D facial animation, largely due to the recent boom in generative models [1, 2, 3, 4, 5, 6, 7, 8] and audio-visual datasets [9, 10, 11, 12, 13]. Recent speech-driven 3D facial animation research is dedicated to generating mouth movements [12, 14, 15, 16, 17, 18, 19], head gestures [20, 21, 19] and emotional facial expressions [22, 23, 24], solely from speech signals. A considerable focus is on achieving synchronization between the animated mouth movements and the corresponding speech. However, the speech-driven animation of an indispensable facial component, eye gaze, has been overlooked by the recent research. Eyes, our “windows to the soul”, serve a crucial role in indicating an individual’s emotions, thoughts, feelings and intentions, thus enabling us to “see someone’s soul” through observing their eyes. In the absence of animated eye gaze, the synthesized 3D talking faces [12, 14, 15, 16, 17, 20, 21, 18, 19, 22, 23, 24] tend to look straight ahead with a static eye pose, resulting in an unnatural and wooden looking facial behavior. Although the recent scene-driven methods [25, 26] achieve impressive eye gaze animation for virtual characters, they require other multi-modal cues besides speech as input, such as scene image, scene context, director script and others.
@@
In this work, we focus on task of generating the expressive 3D eye gaze motion from speech alone, which is in practice challenging for two reasons. 1) Weak Correlation. Eye gaze is one of many non-verbal cues used in communication. It is possible for a speaker to direct his gaze elsewhere while still effectively communicating his message via other verbal means. Mouth strongly correlates with speech, while eye gaze is weakly correlated. As a result, there exists a cross-modal one-to-many mapping between speech signal and eye gaze motion, which is in practice hard to be captured. 2) Data Scarcity. Existing audio-mesh datasets including VOCASET [12], BIWI [13], Speech4Mesh [27], Audio2Mesh [28] and DiffPoseTalk [21] focus on collecting facial motions (mainly mouth movements), but neglect the eye gaze capturing. A 3D dataset consisted of eye gaze motion, head motion and facial motion and their corresponding speech recordings is really scarce.
@@
When we revisit the early research on computer animation, we find that pioneers in this field had proposed a few potential solutions [29, 30, 31, 32, 33], roughly divided into rule-based [29, 30] and learning-based methods [31, 32, 33]. However, both types of methods can only establish a one-to-one mapping between speech signal and eye gaze motion, resulting in deterministic and over-smoothed motion. Additionally, their motion capture systems used to collect eye gaze data in the lab, are costly and time-consuming for constructing large-scale in-the-wild dataset. Ultimately, these methods are mostly customized for specific avatars, making them difficult to integrate with generic 3D face models like FLAME [34].
@@
Figure 1: Given a speech audio, our TalkingEyes can synthesize pluralistic eye gaze motions, eye blinks and head motions and facial motions, and further be used to drive a 3D Gaussian Splatting (3DGS) [35] based head avatar [36].
@@
Secondly, we propose to model the eye gaze motion and head motion across two latent spaces, and subsequently predict them from speech via a temporal autoregressive model. A cascade of a variational autoencoder alongside a vector-quantized variational autoencoder (VAE-VQVAE) is utilized to learn continuous latent space for head motion and discrete latent space for eye gaze motion. This distinction is rooted in the different degrees of motion diversity, caused by the different rotation ranges of eyeballs and head. The head motion is generated and then used as condition in the generation of eye gaze motion, with respect to the synchronous nature of head and eye movements [37, 38, 32]. Furthermore, eye blinks are synthesized based on a statistical analysis on TKED. To create holistic animation, a speech-driven facial animation method [19] is employed to produce mouth and upper-face movements.
@@
  • The cascade of VAE and VQVAE is proposed to model different degrees of diversity in head and eye gaze motions. This facilitates the generation of non-deterministic, diverse and natural eye gaze motions, in contrast with the deterministic generation in existing methods.
@@ II Related Work
The 3D eye gaze animation can be driven by audio, image and video, and is also close related to 3D facial animation. We introduce representative works in these fields, and recommend a comprehensive review [39] of inter-disciplinary research on eye gaze.
@@ II-A Speech-Driven 3D Eye Gaze Animation
It’s a challenging task to generate 3D eye gaze animation from speech, and most existing methods were proposed a decade ago. The methods can be roughly categorized into rule-based [29, 30] and learning-based [31, 32, 33].
Rule-based methods construct the mapping between conversation states and eye gaze patterns by defining animation rules. Masuko and Hoshino [29] designed a set of empirical rules to map the conversation states (e.g. utterance, listening and waiting) to the amount of gaze and gaze duration. Marsella et al. [30] proposed a rule system to sequentially convert the audio and text features to communicative functions, behavior classes, specific behaviors and finally animation.
In contrast, learning-based methods capture the relationship between speech signal and eye movements by training on data. Mariooryad and Busso [31] employed three Dynamic Bayesian Networks to model the coupling between speech, eyebrow and head motion. Le et al. [32] utilized non-linear dynamical canonical correlation analysis to synthesize gaze from head motion and speech features. Jin et al. [33] proposed a LSTM-based machine learning model to predict the direction-of-focus of all the interlocutors in a three-party conversation. All the above learning-based methods can generate head-eye animation from speech, but they only establish a one-to-one mapping between speech feature and eye gaze motion, without the modeling of uncertainty in nonverbal communication.
To reduce uncertainty, some recent scene-driven (not solely speech-driven) methods [25, 26] forced the animated eye gaze to focus on the interlocutors or on the salient points within the scene. Our method only takes the speech as input and achieves the goal of generating plausible eye gaze motion in a speech-centric setting.
@@ II-B Eye Gaze Fitting in Video
Compared with speech, using video to generate eye gaze motion is more straightforward and convenient. The researchers in computer graphics have proposed several 3D face trackers for video. These trackers can reconstruct not only 3D face model but also 3D eye gaze in each frame, by fitting the parameters of both 3D face state and eye gaze state to the image cues. Wang et al. [40] utilized a random forest classifier to extract the iris and pupil pixels in the eye region, and then used them to sequentially infer the most likely state of the 3D eye gaze at each frame in the MAP framework. Wang et al. [41] improved the work [40] by using CNN rather than random forest classifier for the iris and pupil pixels classification. Wen et al. [42] introduced a method for estimating eyeball motions for RGBD inputs by minimizing the differences between a rendered eyeball and a recorded image.
@@ II-C Gaze Estimation from Image
The computer vision community starts very early to estimate gaze from image. The estimated gaze is usually in the form of a unit gaze direction vector in 3D space or a point of gaze (PoG) on 2D plane. Early gaze estimation methods are mainly template-matching-based, appearance-based and feature-based, while recent methods are mostly based on deep learning [43]. Zhang et al. [44] introduced the first CNN-based method for gaze estimation, which jointly utilizes eye images and head poses to predict gaze directions. Accounting for the potential asymmetry between human eyes, Cheng et al. [45] proposed to firstly predict the 3D gaze direction for each eye individually and then employ a reliability evaluation network to adaptively adjust the weighting strategy throughout the optimization. Park et al.[46] introduced a novel graphical representation of 3D gaze direction, which serves as intermediate supervision to improve the estimation accuracy. Owing to individual differences, some methods [47, 48] employed a few-shot approach to achieve user-specific gaze adaptation. To address the challenge of occluded eyes, Kellnhofer et al. [49] utilized visible head features to predict gaze direction, even when the eyes are completely obscured, and indicated the prediction’s limited accuracy by outputting a higher uncertainty value accordingly.
@@ II-D Speech-Driven 3D Facial Animation
Methods in this field are roughly categorized into linguistics-based [50, 51, 52, 53] and learning-based [12, 14, 15, 16, 17, 21, 19, 22, 23, 24]. Linguistics-based methods typically establish a set of mapping rules between phonemes and visemes, thus providing explicit control over the animation. JALI [50] is a representative recent linguistics-based method, which utilizes two anatomical actions (jaw and lip) to animate a 3D facial rig. As a lot of manual effort is required for tuning animation parameters, a wide variety of data-driven methods have been proposed as an alternative. The generative models are utilized in these methods to directly learn the mapping from speech features to 3D facial motion, which include CNN [12, 14], Transformer [15, 16, 22, 23, 19] and Diffusion Model [17, 21, 24]. VOCASET [12] and BIWI [13] are two widely used 3D audio-visual datasets to train the audio-to-mesh regression networks in the learning-based methods.
Considering that Learn2Talk [19] has achieved impressive lip-sync facial animation, we choose it as the 3D facial motion generator in our proposed method.
@@
VOCASET[12] - - 2019 480 12 0.48 3D Scanning /
BIWI[13] - - 2010 1109 14 1.43 3D Scanning /
Speech4Mesh[27] - - 2023 2000 / / Generated from video MEAD[10], VoxCeleb2[9]
Audio2Mesh[28] - - 2024 >1M 6112 / Generated from video VoxCeleb2[9]
DiffPoseTalk[21] - 2024 1052 588 26.5 Generated from video TFHP[21], HDTF[11]
Our TKED 2025 5982 669 14.0 Generated from video VoxCeleb2[9], HDTF [11]
@@ III-A Dataset Description
The way of utilizing motion capture systems or eye tracking devices to capture eye gaze and microphone to record audio in a laboratory setting is limited in its ability to collect large-scale and diverse data. Considering that 2D audio-visual datasets are more accessible and offer broader coverage, it’s possible to construct the audio-gaze dataset by reconstructing both 3D face mesh and 3D eye gaze from video. Motivated by this, we propose TKED which is obtained by performing 3D face reconstruction and 3D eye gaze fitting on the VoxCeleb2 [9] and HDTF [11] datasets.
@@
TKED consists of 5,982 videos featuring 669 subjects, amounting to approximately 14 hours of footage in total. Out of the 5,982 videos, 5,912 are from VoxCeleb2, with a total duration of about 12 hours, while the remaining 70 videos are from HDTF, totaling about 2 hours. To ensure consistency, the videos from HDTF have been resampled to match the 25 fps in VoxCeleb2. We have conducted 3D face reconstruction and 3D eye gaze fitting on each frame of these videos to obtain the pseudo ground truth FLAME parameters [34], which include two eyeballs’ rotation, head rotation and facial motion parameters. Fig. 2 shows four sequences of the reconstructed 3D FLAME model. Along with the audio stored in the original videos, TKED is first dataset in the research community to simultaneously contain the three types of resources: 3D talking eyes, 3D talking head and 3D talking faces. The comprehensive comparison between TKED and existing audio-mesh datasets [12, 13, 27, 28, 21] is presented in Tab. I. The proportion of training, validation and testing data in TKED is 8:1:1.

### v2

@@ Abstract
Although significant progress has been made in the field of speech-driven 3D facial animation recently, the speech-driven animation of an indispensable facial component, eye gaze, has been overlooked by recent research. This is primarily due to the weak correlation between speech and eye gaze, as well as the scarcity of audio-gaze data, making it very challenging to generate 3D eye gaze motion from speech alone. In this paper, we propose a novel data-driven method which can generate diverse 3D eye gaze motions in harmony with the speech. To achieve this, we firstly construct an audio-gaze dataset that contains about 14 hours of audio-mesh sequences featuring high-quality eye gaze motion, head motion and facial motion simultaneously. The motion data is acquired by performing lightweight eye gaze fitting and face reconstruction on videos from existing audio-visual datasets. We then tailor a novel speech-to-motion translation framework in which the head motions and eye gaze motions are jointly generated from speech but are modeled in two separate latent spaces. This design stems from the physiological knowledge that the rotation range of eyeballs is less than that of head. Through mapping the speech embedding into the two latent spaces, the difficulty in modeling the weak correlation between speech and non-verbal motion is thus attenuated. Finally, our TalkingEyes, integrated with a speech-driven 3D facial motion generator, can synthesize eye gaze motion, eye blinks, head motion and facial motion collectively from speech. Qualitative and quantitative evaluations, along with a perceptual user study, demonstrate the superiority of the proposed method in generating diverse and natural 3D eye gaze motions from speech. The project page of this paper is: https://lkjkjoiuiu.github.io/TalkingEyes_Home/.
@@ I Introduction
The computer graphics and vision community has witnessed a significant progress in the field of speech-driven 3D facial animation, largely due to the recent boom in generative models [vae2014, VQVAE2017, GANs2014, StyleGAN2019, Attention2017, DDPM2020, StableDiffusion, DiffSFSR2024] and audio-visual datasets [chung2018voxceleb2, wang2020mead, zhang2021flow, VOCA2019, BIWI2010]. Recent speech-driven 3D facial animation research is dedicated to generating mouth movements [VOCA2019, MeshTalk2021, FaceFormer2022, CodeTalker2023, FaceDiffuser2023, DNPMICME2024, Learn2Talk2025], head gestures [3DTalkingFaceHead2023, DiffPoseTalkTOG2024, Learn2Talk2025] and emotional facial expressions [EmoTalk2023, EMOTE2023, Media2FaceSIG2024], solely from speech signals. A considerable focus is on achieving synchronization between the animated mouth movements and the corresponding speech. However, the speech-driven animation of an indispensable facial component, eye gaze, has been overlooked by the recent research. Eyes, our “windows to the soul”, serve a crucial role in indicating an individual’s emotions, thoughts, feelings and intentions, thus enabling us to “see someone’s soul” through observing their eyes. In the absence of animated eye gaze, the synthesized 3D talking faces [VOCA2019, MeshTalk2021, FaceFormer2022, CodeTalker2023, FaceDiffuser2023, 3DTalkingFaceHead2023, DiffPoseTalkTOG2024, DNPMICME2024, Learn2Talk2025, EmoTalk2023, EMOTE2023, Media2FaceSIG2024] tend to look straight ahead with a static eye pose, resulting in an unnatural and wooden looking facial behavior. Although the recent scene-driven methods [SaliencyGazeTVCG2024, S3TOG2024] achieve impressive eye gaze animation for virtual characters, they require other multi-modal cues besides speech as input, such as scene image, scene context, director script and others.
@@
In this work, we focus on task of generating the expressive 3D eye gaze motion from speech alone, which is in practice challenging for two reasons. 1) Weak Correlation. Eye gaze is one of many non-verbal cues used in communication. It is possible for a speaker to direct his gaze elsewhere while still effectively communicating his message via other verbal means. Mouth strongly correlates with speech, while eye gaze is weakly correlated. As a result, there exists a cross-modal one-to-many mapping between speech signal and eye gaze motion, which is in practice hard to be captured. 2) Data Scarcity. Existing audio-mesh datasets including VOCASET [VOCA2019], BIWI [BIWI2010], Speech4Mesh [he2023speech4mesh], Audio2Mesh [yang2024probabilistic] and DiffPoseTalk [DiffPoseTalkTOG2024] focus on collecting facial motions (mainly mouth movements), but neglect the eye gaze capturing. A 3D dataset consisted of eye gaze motion, head motion and facial motion and their corresponding speech recordings is really scarce.
@@
When we revisit the early research on computer animation, we find that pioneers in this field had proposed a few potential solutions [HeadEyeAnimation2007, VirtualCharacterSCA2013, ConversationalAgents2012, LiveSpeehEyeTVCG2012, EyeMotion2019], roughly divided into rule-based [HeadEyeAnimation2007, VirtualCharacterSCA2013] and learning-based methods [ConversationalAgents2012, LiveSpeehEyeTVCG2012, EyeMotion2019]. However, both types of methods can only establish a one-to-one mapping between speech signal and eye gaze motion, resulting in deterministic and over-smoothed motion. Additionally, their motion capture systems used to collect eye gaze data in the lab, are costly and time-consuming for constructing large-scale in-the-wild dataset. Ultimately, these methods are mostly customized for specific avatars, making them difficult to integrate with generic 3D face models like FLAME [FLAME2017].
@@
Figure 1: Given a speech audio, our TalkingEyes can synthesize pluralistic eye gaze motions, eye blinks and head motions and facial motions, and further be used to drive a 3D Gaussian Splatting (3DGS) [3DGSTOG2023] based head avatar [GaussianAvatarsCVPR2024].
@@
Secondly, we propose to model the eye gaze motion and head motion across two latent spaces, and subsequently predict them from speech via a temporal autoregressive model. We employ a variational autoencoder (VAE) to learn continuous latent space for head motion and a vector-quantized variational autoencoder (VQVAE) to learn discrete latent space for eye gaze motion. This distinction is rooted in the different degrees of motion diversity, caused by the different rotation ranges of eyeballs and head. The head motion is firstly generated from speech and then used as a condition in the autoregressive generation of eye gaze motion, with respect to the synchronous nature of head and eye movements [Ophthal1972, Warabi1977, LiveSpeehEyeTVCG2012]. Qualitative and quantitative evaluations, along with a perceptual user study, demonstrate the superiority of the proposed TalkingEyes in generating diverse and natural 3D eye gaze motions from speech.
To create holistic animation, the eye blinks are synthesized based on a statistical analysis on TKED, and the mouth and upper-face movements are generated by a speech-driven facial animation method [Learn2Talk2025].
@@
  • VAE and VQVAE are proposed to model different degrees of diversity in head and eye gaze motions. This facilitates the generation of non-deterministic, diverse and natural eye gaze motions, in contrast with the deterministic generation in existing methods.
@@ II Related Work
The 3D eye gaze animation can be driven by audio, image and video, and is also close related to 3D facial animation. We introduce representative works in these fields, and recommend a comprehensive review [EyeGazeReview2015] of inter-disciplinary research on eye gaze.
@@ II-A Speech-Driven 3D Eye Gaze Animation
It’s a challenging task to generate 3D eye gaze animation from speech, and most existing methods were proposed a decade ago. The methods can be roughly categorized into rule-based [HeadEyeAnimation2007, VirtualCharacterSCA2013] and learning-based [ConversationalAgents2012, LiveSpeehEyeTVCG2012, EyeMotion2019].
Rule-based methods construct the mapping between conversation states and eye gaze patterns by defining animation rules. Masuko and Hoshino [HeadEyeAnimation2007] designed a set of empirical rules to map the conversation states (e.g. utterance, listening and waiting) to the amount of gaze and gaze duration. Marsella et al. [VirtualCharacterSCA2013] proposed a rule system to sequentially convert the audio and text features to communicative functions, behavior classes, specific behaviors and finally animation.
In contrast, learning-based methods capture the relationship between speech signal and eye movements by training on data. Mariooryad and Busso [ConversationalAgents2012] employed three Dynamic Bayesian Networks to model the coupling between speech, eyebrow and head motion. Le et al. [LiveSpeehEyeTVCG2012] utilized non-linear dynamical canonical correlation analysis to synthesize gaze from head motion and speech features. Jin et al. [EyeMotion2019] proposed a LSTM-based machine learning model to predict the direction-of-focus of all the interlocutors in a three-party conversation. All the above learning-based methods can generate head-eye animation from speech, but they only establish a one-to-one mapping between speech feature and eye gaze motion, without the modeling of uncertainty in nonverbal communication.
To reduce uncertainty, some recent scene-driven (not solely speech-driven) methods [SaliencyGazeTVCG2024, S3TOG2024, migCanalesJJ23] forced the animated eye gaze to focus on the interlocutors or on the salient points within the scene. Our method only takes the speech as input and achieves the goal of generating plausible eye gaze motion in a speech-centric setting.
@@ II-B Eye Gaze Fitting in Video
Compared with speech, using video to generate eye gaze motion is more straightforward and convenient. The researchers in computer graphics have proposed several 3D face trackers for video. Extending 3D face reconstruction methods [Deep3DFace2019, DECA2021, DCT2024, FaceRefiner2024, DNPMICME2024], these trackers jointly fit parameters for both face and eye gaze states to image cues, enabling them to reconstruct a 3D face model along with 3D eye gaze for each frame. Wang et al. [wang2016realtime] utilized a random forest classifier to extract the iris and pupil pixels in the eye region, and then used them to sequentially infer the most likely state of the 3D eye gaze at each frame in the MAP framework. Wang et al. [wang2019realtime] improved the work [wang2016realtime] by using CNN rather than random forest classifier for the iris and pupil pixels classification. Wen et al. [wen2017real] introduced a method for estimating eyeball motions for RGBD inputs by minimizing the differences between a rendered eyeball and a recorded image.
@@ II-C Gaze Estimation from Image
The computer vision community starts very early to estimate gaze from image. The estimated gaze is usually in the form of a unit gaze direction vector in 3D space or a point of gaze (PoG) on 2D plane. Early gaze estimation methods are mainly template-matching-based, appearance-based and feature-based, while recent methods are mostly based on deep learning [GazeEstimationSurvey2024]. Zhang et al. [zhang2015appearance] introduced the first CNN-based method for gaze estimation, which jointly utilizes eye images and head poses to predict gaze directions. Accounting for the potential asymmetry between human eyes, Cheng et al. [cheng2020gaze] proposed to firstly predict the 3D gaze direction for each eye individually and then employ a reliability evaluation network to adaptively adjust the weighting strategy throughout the optimization. Park et al.[park2018deep] introduced a novel graphical representation of 3D gaze direction, which serves as intermediate supervision to improve the estimation accuracy. Owing to individual differences, some methods [park2019few, yu2019improving] employed a few-shot approach to achieve user-specific gaze adaptation. To address the challenge of occluded eyes, Kellnhofer et al. [kellnhofer2019gaze360] utilized visible head features to predict gaze direction, even when the eyes are completely obscured, and indicated the prediction’s limited accuracy by outputting a higher uncertainty value accordingly.
@@ II-D Speech-Driven 3D Facial Animation
Methods in this field are roughly categorized into linguistics-based [JALI2016, AnimatedSpeech2001, DynamicUnits2012, LipSyncGames2013] and learning-based [VOCA2019, MeshTalk2021, FaceFormer2022, CodeTalker2023, FaceDiffuser2023, DiffPoseTalkTOG2024, Learn2Talk2025, EmoTalk2023, EMOTE2023, Media2FaceSIG2024]. Linguistics-based methods typically establish a set of mapping rules between phonemes and visemes, thus providing explicit control over the animation. JALI [JALI2016] is a representative recent linguistics-based method, which utilizes two anatomical actions (jaw and lip) to animate a 3D facial rig. As a lot of manual effort is required for tuning animation parameters, a wide variety of data-driven methods have been proposed as an alternative. The generative models are utilized in these methods to directly learn the mapping from speech features to 3D facial motion, which include CNN [VOCA2019, MeshTalk2021], Transformer [FaceFormer2022, CodeTalker2023, EmoTalk2023, EMOTE2023, Learn2Talk2025] and Diffusion Model [FaceDiffuser2023, DiffPoseTalkTOG2024, Media2FaceSIG2024]. VOCASET [VOCA2019] and BIWI [BIWI2010] are two widely used 3D audio-visual datasets to train the audio-to-mesh regression networks in the learning-based methods.
Considering that Learn2Talk [Learn2Talk2025] has achieved impressive lip-sync facial animation, we choose it as the 3D facial motion generator in our proposed method.
@@
VOCASET[VOCA2019] - - 2019 480 12 0.48 3D Scanning /
BIWI[BIWI2010] - - 2010 1109 14 1.43 3D Scanning /
Speech4Mesh[he2023speech4mesh] - - 2023 2000 / / Generated from video MEAD[wang2020mead], VoxCeleb2[chung2018voxceleb2]
Audio2Mesh[yang2024probabilistic] - - 2024 >1M 6112 / Generated from video VoxCeleb2[chung2018voxceleb2]
DiffPoseTalk[DiffPoseTalkTOG2024] - 2024 1052 588 26.5 Generated from video TFHP[DiffPoseTalkTOG2024], HDTF[zhang2021flow]
Our TKED 2025 5982 669 14.0 Generated from video VoxCeleb2[chung2018voxceleb2], HDTF [zhang2021flow]
@@ III-A Dataset Description
The way of utilizing motion capture systems or eye tracking devices to capture eye gaze and microphone to record audio in a laboratory setting is limited in its ability to collect large-scale and diverse data. Considering that 2D audio-visual datasets are more accessible and offer broader coverage, it’s possible to construct the audio-gaze dataset by reconstructing both 3D face mesh and 3D eye gaze from video. Motivated by this, we propose TKED which is obtained by performing 3D face reconstruction and 3D eye gaze fitting on the VoxCeleb2 [chung2018voxceleb2] and HDTF [zhang2021flow] datasets.
@@
TKED consists of 5,982 videos featuring 669 subjects, amounting to approximately 14 hours of footage in total. Out of the 5,982 videos, 5,912 are from VoxCeleb2, with a total duration of about 12 hours, while the remaining 70 videos are from HDTF, totaling about 2 hours. To ensure consistency, the videos from HDTF have been resampled to match the 25 fps in VoxCeleb2. We have conducted 3D face reconstruction and 3D eye gaze fitting on each frame of these videos to obtain the pseudo ground truth FLAME parameters [FLAME2017], which include two eyeballs’ rotation, head rotation and facial motion parameters. Fig. 2 shows four sequences of the reconstructed 3D FLAME model. Along with the audio stored in the original videos, TKED is first dataset in the research community to simultaneously contain the three types of resources: 3D talking eyes, 3D talking head and 3D talking faces. The comprehensive comparison between TKED and existing audio-mesh datasets [VOCA2019, BIWI2010, he2023speech4mesh, yang2024probabilistic, DiffPoseTalkTOG2024] is presented in Tab. I. The training, validation and testing data in TKED are split in an 8:1:1 ratio based on the number of video clips.